# Measurement, and Mitigation

### v1

@@ 6.1 Training-time interventions
Data cleaning plays a crucial role in mitigating memorization by limiting an LLM’s exposure to overrepresented sequences. As demonstrated by Carlini et al. (2021), Lee et al. (2022), and Kandpal et al. (2022), de-duplication minimizes the over-representation of repeated sequences. Removing these duplicates serves as a form of implicit regularization, reducing the likelihood that the LLM overfits.
@@
From an optimization perspective, LLMs can minimize loss on rare training examples by memorizing. This makes PII-scrubbing effective: by removing these sequences during training, memorization is minimized (Carlini et al., 2021; Jagielski et al., 2023).
@@
Differential privacy (DP) provides robust protection against membership inference and extraction attacks, even in the presence of adversaries with auxiliary knowledge (Dwork, 2006; Dwork et al., 2014; Shokri et al., 2017; Carlini et al., 2021). Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016) guarantees privacy by computing per-example gradients, clipping them to a fixed norm to bound individual influence, and injecting calibrated Gaussian noise into the aggregated updates. A privacy accountant tracks cumulative privacy loss to enforce the global ($\epsilon$, $\delta$) budget.
@@
Li et al. (2021) demonstrated that LLMs can act as strong DP learners. Their evaluation reveals that pretrained LLMs are highly resilient to the noise introduced by DP-SGD, particularly when fine-tuned on top of general-purpose representations. DP-trained models approach the performance of non-private baselines, while offering provable protection against memorization. Empirical validation through canary insertion and MIAs confirm the effectiveness of DP in preventing leakage of sensitive sequences. To further reduce memorization, Zhao et al. (2022a) introduced Confidentially Redacted Training (CRT), a method that combines de-duplication and redaction with DP-SGD to train LLMs while avoiding the retention of sensitive content. CRT builds on DP concepts to introduce randomized training interventions that prevent unintended memorization. The authors show that CRT, when combined with DP-SGD, provides strong privacy protections without compromising model performance, as indicated by competitive perplexity scores.
@@
However, the use of DP in LLMs often presents trade-offs between privacy, utility, and computational efficiency. Strict privacy budgets often lead to performance drops, and significant computational overhead makes it challenging to scale to LLMs. To address these challenges, recent work has examined parameter-efficient fine-tuning (PEFT) strategies for training under DP. PEFT approaches, such as LoRA (Hu et al., 2022), allow finetuning of models with addition of few parameters relative to base models via low-rank decompositions of additional trainable parameters. It is hypothesized that finetuning PEFT models with DP mitigates losses of utility due to requiring less noise for privatization due to training with fewer parameters (Liu et al., 2024b; Yu et al., 2022, 2021). Ma et al. (2024) explored the integration of adapter-based fine-tuning methods into the DP framework. Their findings reveal that DP PEFT reduces memorization and achieves comparable or superior downstream task performance relative to full-model DP-SGD, under strict privacy budgets. Ma et al. (2024) hypothesized that the limited parameter updates in PEFT may concentrate DP noise in a narrow subset of the model, potentially weakening privacy protection. While most DP techniques for LLMs target record-level privacy, many applications require user-level guarantees. Chua et al. (2024) addressed this gap by proposing several techniques to enforce DP at user-level. To evaluate effectiveness, Chua et al. (2024) used canary insertion attacks, embedding unique sequences into user data to test for memorization. Results show that user-level DP reduces memorization, with much lower canary extraction rates than record-level DP.
@@
DP in LLMs continue to present trade-offs between privacy, utility, and computational efficiency. Strict privacy budgets often lead to performance drops, and significant computational overhead makes it challenging to scale to LLMs. Thus, the adoption of DP for mitigating memorization in LLMs remains limited.
@@ 6.2 Post-training-time interventions
Machine unlearning aims to remove the influence of certain training examples so that the model’s behavior is as if those examples were never seen (Cooper et al., 2024; Liu, 2024).  Yao et al. (2024b) evaluates various model-editing unlearning strategies (e.g., gradient ascent (Golatkar et al., 2020; Jia et al., 2023; Jang et al., 2022), negative re-labeling (Golatkar et al., 2020), adversarial sampling (Cha et al., 2024)) demonstrating that approximate unlearning methods can be over $10^{5}\times$ more computationally efficient than retraining a model from scratch. However, unlike DP, there is no formal guarantee, thereby leaving a risk that memorization persists.
ParaPO, introduced by Chen et al. (2025), is a post-training strategy to decrease memorization of data memorized in the pretraining corpus. ParaPO decreases unintended memorization by first assessing which data from the pretraining corpus has been memorized (by searching for sequences that, when a prefix is used to prompt an LM, an LM decodes a near exact match to the suffix of the sequence), and then creating preference pairs by using a separate LLM to summarize the memorized datapoint. The model is then post-trained via DPO (Rafailov et al., 2024) on pairs of (memorized sequence, summarization), with the summarization marked as the preferred response. Chen et al. (2025) show that this approach decreases unintended memorization while preserving verbatim recall of desired sequences (e.g. direct quotations) However, the authors note that this approach slightly decreases utility on math, knowledge and reasoning benchmarks.
Machine unlearning and ParaPO constitute post-training interventions for reducing memorization. ParaPO results in slight degradation of utility, raising the question of what extent memorization is required for utility and generalization. Additionally, a future research direction may be to what extent this tradeoff can be tuned via incorporation of “memorization” reward models that are optimized in conjunction with utility-oriented reward models, in the context of RLHF or RLVR.
@@ 6.3 Inference-time interventions
MemFree decoding is a strategy introduced by Ippolito et al. (2023) to filter out memorized sequences during generation. This approach contrasts with decoding by introducing a privacy filter into the generation loop applied post hoc at generation and acts as a wrapper around any pre-trained model. Ippolito et al. (2023) leveraged a bloom filter to represent the set of all n-grams in the training set to identify memorized content in real time. However, near-identical n-grams are undetectable as minor modifications to a token sequence evade detection. Another limitation is that access to the LLM’s n-gram training data is required.
@@
Activation steering addresses memorization in LLMs by manipulating internal activations during inference (Turner et al., 2023). By injecting targeted perturbations into an LLM’s hidden states, activation steering can suppress or redirect the generation of memorized sequences while preserving overall model reasoning capabilities (Suri et al., 2025). Suri et al. (2025) demonstrated that using a sparse autoencoder to identify activation patterns linked to memorized passages enables the construction of steering directions that reduced memorization by up to 60%, with minimal performance degradation. However, they note that steering effectiveness depends heavily on precise layer selection and steering strength.
@@
Understanding where memorization arises in LLMs is crucial to designing effective steering interventions. Stoehr et al. (2024) conducted a detailed mechanistic analysis of paragraph-length memorized sequences and found that memorization is often localized to specific LLM components. They showed that fine-tuning only a subset of the high-gradient weights was sufficient to erase memorized passages. Additionally, they identified a single attention head in an early layer that reliably activates in response to rare token combinations seen during training, triggering verbatim recall.
@@
Building on localization, Chang et al. (2023) evaluated multiple methods for localizing memorized content by introducing two diagnostic benchmarks: injection, where known memorized sequences are inserted via fine-tuning a small set of weights, and deletion, which tests whether removing specific neurons erases a naturally memorized output. They compared several techniques, including activation-based heuristics (Geva et al., 2022), integrated gradients (Dai et al., 2021), and pruning-based methods like Slimming and Hard Concrete (Chang et al., 2023). Their results showed that pruning-based approaches were most effective: for example, Hard Concrete was able to identify fewer than 0.5% of neurons whose removal led to a  60% drop in memorization accuracy. Both Chang et al. (2023) and Stoehr et al. (2024) highlighted a key challenge: the neurons involved in one memory often contribute to others, complicating memory-specific interventions due to the risk of collateral forgetting.
@@ 6.4 Connections to other undesirable behaviors
It remains open whether mitigating memorization also mitigates other undesirable behaviors of LLMs, such as hallucination, racial bias, and toxic content. For instance, does reducing memorization reduce hallucination on concepts related to memorized data? Is it possible that reducing the memorization of datapoints that include harmful racial stereotypes reduces the rates at which large models produce biased or toxic content? The authors to date have not seen cross-cutting analysis examining whether reducing memorization can reduce other harmful behaviors in-tandem.
@@
Open Questions 1. How can LLMs be designed or trained for selective memorization? 2. Shall we scale up DP training, and at what cost and utility trade-off? 3. How can DP improve parameter-efficient fine-tuning to reduce memorization? 4. How can we optimize activation steering for memorization removal? 5. How can post-training methods reduce memorization while preserving utility? 6. Can memorization mitigation methods be incorporated online during post-training, for instance, by using memorization-detecting reward models during RLHF? 7. When does memorization contribute to versus harm utility/generalization? 8. Does mitigating memorization affect other undesirable behaviors, such as hallucination or generating toxic stereotypes?
@@ 7 Privacy & legal risks of memorization
Memorization in LLMs poses serious security and privacy risks: memorized text can be directly leaked, undermining data confidentiality (Smith et al., 2023). We discuss three major impact areas caused by memorization: personal data leakage, exposure of copyrighted or proprietary content, and broader legal and public consequences.
@@
Personal data leakage. As explored in Li et al. (2024), Carlini et al. (2021), Jagielski et al. (2023), and Zhou et al. (2024), a significant risk for LLMs is leaking sensitive personal information. In Panda et al. (2024), researchers demonstrated a targeted neural phishing attack that achieved up to a 50% success rate in tricking a model into revealing PII during fine-tuning when injecting poisons during pretraining. Even when training data is intended to remain private, malicious attackers with the correct model access level can siphon memorized secrets (Panda et al., 2024). If an LLM reveals any PII, this could result in identity theft or other harms. Furthermore, for industries such as healthcare or customer service, where customers expect confidentiality, this would place model operators in the crosshairs of current privacy regulations.
@@
Copyright or proprietary content. Beyond PII, memorization in LLMs may also cause the reproduction of copyrighted or proprietary material such as books, code, etc. This turns the LLM into a potential conduit for the unauthorized distribution of protected content. As observed by Carlini et al. (2023), models may produce verbatim memorization from their training data. While we can only explicitly confirm what training data is used for open-source LLMs (Gao et al., 2020; Soldaini et al., 2024), there do exist copyrighted books in EleutherAI’s PILE dataset, specifically, copyrighted books in its Books3 section. This is observed in Henderson et al. (2023) and Zhang et al. (2022), where open models can regurgitate the first few pages of the Harry Potter books and “Oh the places you’ll go!” by Dr. Seuss verbatim, raising questions regarding copyright.
@@
Legal landscape. There are several ongoing cases that will shape the legal landscape of LLMs. The New York Times v. Microsoft Corp. lawsuit (NYT-MC, 2023) represents a pivotal legal challenge to generative AI and copyright, accompanied by similar cases, including Chabon v. OpenAI (Chabon-OpenAI, 2023) where authors claiming copyright infringement in model training, and Doe v. GitHub (DOE-GitHub, 2022) where plaintiffs allege verbatim code reproduction in violation of the DMCA. The Times claimed that OpenAI’s model not only trained on its copyrighted articles without proper license but also that ChatGPT could regurgitate substantial passages of those articles on request. In NYT-MC (2023), the publishers argue this behavior amounts to mass copyright infringement since the model’s outputs mimic and even compete with the Times’ content. While researchers such as  Freeman et al. (2024) may provide their analysis regarding the veracity of lawsuit (NYT-MC, 2023)’s copyright infringement claims, such high-profile legal action underscores how memorization is no longer only a technical but a business and public policy problem.

### v2

@@ 6.1 Training-Time Interventions
Data cleaning plays a crucial role in mitigating memorization by limiting an LLM’s exposure to overrepresented sequences. As demonstrated by [carlini2021extractingtrainingdatalarge, lee2022deduplicatingtrainingdatamakes, kandpal2022deduplicatingtrainingdatamitigates], de-duplication minimizes the over-representation of repeated sequences. Removing these duplicates serves as a form of implicit regularization, reducing the likelihood that the LLM overfits to this redundant data.
@@
From an optimization perspective, LLMs can minimize loss on rare training examples by memorizing. This makes PII-scrubbing effective: by removing these sequences during training, memorization is minimized [carlini2021extractingtrainingdatalarge, jagielski2023measuring].
The necessity of proactive data sanitization is underscored by findings from Li, et al.[li2024llmpbeassessingdataprivacy], who demonstrated that even state-of-the-art models can leak a significant portion of held-out PII when subjected to adversarial “true-prefix” prompts. Effective mitigation relies on systematic PII scrubbing, a process guided by frameworks such as the taxonomy developed by Lukas, et al.[lukas2023analyzing], which categorizes PII types (e.g., CARDINAL, DATE, PERSON) to enable both rule-based and machine-learning-based removal of sensitive tokens. The application of these scrubbing guidelines was shown to dramatically reduce such leakage, confirming the efficacy of targeted data cleaning as a primary defense mechanism. This principle of targeted mitigation extends beyond PII to other forms of protected content, such as copyrighted material. For instance, Vyas, et al. [vyas2023provable] introduced a method for training models to mitigate near-exact copying of copyrighted excerpts, providing provable certificates by extending differential privacy concepts to a bounded memorization budget.
@@
Differential privacy (DP) [brown2022does] provides robust protection against membership inference and extraction attacks, even in the presence of adversaries with auxiliary knowledge [dwork2006differential, dwork2014algorithmic, shokri2017membership, carlini2021extractingtrainingdatalarge]. Differentially Private Stochastic Gradient Descent (DP-SGD) [abadi2016deep] guarantees privacy by computing per-example gradients, clipping them to a fixed norm to bound individual influence, and injecting calibrated Gaussian noise into the aggregated updates. A privacy accountant tracks cumulative privacy loss to enforce the global ($\epsilon$, $\delta$) budget.
@@
Li, et al. [li2021large] demonstrated that LLMs can act as strong DP learners. Their evaluation reveals that pretrained LLMs are highly resilient to the noise introduced by DP-SGD, particularly when fine-tuned on top of general-purpose representations. DP-trained models approach the performance of non-private baselines, while offering provable protection against memorization. Empirical validation through canary insertion and MIAs confirm the effectiveness of DP in preventing leakage of sensitive sequences. To further reduce memorization, Zhao, et al. [zhao2022provably] introduced Confidentially Redacted Training (CRT), a method that combines de-duplication and redaction with DP-SGD to train LLMs while avoiding the retention of sensitive content. CRT builds on DP concepts to introduce randomized training interventions that prevent unintended memorization. The authors show that CRT, when combined with DP-SGD, provides strong privacy protections without compromising model performance, as indicated by competitive perplexity scores. However, integrating DP with pretraining memorization is highly complex at scale. As the provable guarantees of DP rely on bounding the “group size,” the total number of times an individual’s data appears, the vast, web-scale nature of pretraining corpora, coupled with the scale duplication and near-duplicates, make reliably defining this group size extremely challenging. Consequently, the interpretation and strength of DP guarantees against pretraining-time memorization often remain unclear due to the underlying data redundancy.
@@
Furthermore, the use of DP in LLMs often presents trade-offs among privacy, utility, and computational efficiency. Strict privacy budgets often lead to performance drops, and significant computational overhead makes it challenging to scale to LLMs. To address these challenges, recent work has examined PEFT strategies for training under DP. PEFT approaches, such as LoRA [hu2022lora], allow fine-tuning of models with addition of few parameters relative to base models via low-rank decompositions of additional trainable parameters. It is hypothesized that fine-tuning PEFT models with DP mitigates losses of utility due to requiring less noise for privatization due to training with fewer parameters [liu2024differentially, yu2022differentiallyprivatefinetuninglanguage, yu2021largescaleprivatelearning]. Ma, et al. [ma2024efficient] explored the integration of adapter-based fine-tuning methods into the DP framework. Their findings reveal that DP PEFT reduces memorization and achieves comparable or superior downstream task performance relative to full-model DP-SGD, under strict privacy budgets. Ma, et al. [ma2024efficient] hypothesized that the limited parameter updates in PEFT may concentrate DP noise in a narrow subset of the model, potentially weakening privacy protection. While most DP techniques for LLMs target record-level privacy, many applications require user-level guarantees. Chua, et al. [chua2024mind] addressed this gap by proposing several techniques to enforce DP at user-level. To evaluate effectiveness, Chua, et al. [chua2024mind] used canary insertion attacks, embedding unique sequences into user data to test for memorization. Results show that user-level DP reduces memorization, with much lower canary extraction rates than record-level DP.
@@
DP in LLMs faces an efficiency-utility trade-off: strict privacy budgets degrade performance, while scaling DP methods requires significant computational overhead. Thus, the adoption of DP for mitigating memorization in LLMs remains limited.
@@ 6.2 Post-Training-Time Interventions
Machine unlearning aims to remove the influence of certain training examples so that the model’s behavior is as if those examples were never seen [cooper2024machine, liu2024unlearning]. Yao, et al.  [yao2024machine] evaluates various model-editing unlearning strategies (e.g., gradient ascent [golatkar2020eternal, jia2023model, jang2022knowledge], negative re-labeling [golatkar2020eternal], adversarial sampling [cha2024learning]) demonstrating that approximate unlearning methods can be over $10^{5}\times$ more computationally efficient than retraining a model from scratch. However, unlike DP, there is no formal guarantee, thereby leaving a risk that memorization persists.
ParaPO, introduced by [chen2025parapoaligninglanguagemodels], is a post-training strategy to decrease memorization of data memorized in the pretraining corpus. ParaPO decreases unintended memorization by first assessing which data from the pretraining corpus has been memorized (by searching for sequences that, when a prefix is used to prompt an LM, an LM decodes a near exact match to the suffix of the sequence), and then creating preference pairs by using a separate LLM to summarize the memorized datapoint. The model is then post-trained via Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage] on pairs of (memorized sequence, summarization), with the summarization marked as the preferred response. Chen, et al. [chen2025parapoaligninglanguagemodels] show that this approach decreases unintended memorization while preserving verbatim recall of desired sequences (e.g. direct quotations) However, the authors note that this approach slightly decreases utility on math, knowledge and reasoning benchmarks.
Machine unlearning and ParaPO constitute post-training interventions for reducing memorization. ParaPO results in slight degradation of utility, raising the question of what extent memorization is required for utility and generalization. A future research direction may be to what extent this tradeoff can be tuned via incorporation of “memorization” reward models that are optimized in conjunction with utility-oriented reward models, leveraging RLHF or RLVR.
Model alignment processes like instruction tuning and RLHF are primarily designed to make models safer and more helpful, but they have a significant consequential effect on memorization. As the work by Nasr et al. [nasr2023scalableextractiontrainingdata] details, these alignment procedures reduce the likelihood that a model will directly reproduce memorized training data when queried with benign or prefix-style prompts. Thus, alignment makes memorized content harder to extract under normal interaction settings. However, it does not eliminate memorization altogether: with divergence-based prompting or targeted fine-tuning, aligned models can still be induced to reproduce sensitive or verbatim sequences from pretraining data. This finding highlights that alignment primarily limits the accessibility of memorized data rather than its existence, implying that effective mitigation must combine alignment with broader privacy-preserving training or post-hoc filtering strategies. Therefore, alignment makes memorization harder to extract but does not make it harder to memorize (or prevent the memory from existing). This is a crucial distinction between achieving safety/utility and achieving true privacy/unlearning.
@@ 6.3 Inference-Time Interventions
MemFree decoding is a strategy introduced by [ippolito-etal-2023-preventing] to filter out memorized sequences during generation. This approach contrasts with decoding by introducing a privacy filter into the generation loop applied post hoc during generation and acts as a wrapper around any pre-trained model. Ippoloto, et al. [ippolito-etal-2023-preventing] leveraged a bloom filter to represent the set of all n-grams in the training set to identify memorized content in real time. However, near-identical n-grams are undetectable as minor modifications to a token sequence evade detection. Another limitation is that access to the LLM’s n-gram training data is required.
TokenSwap [prashanttokenswap] introduces a lightweight, token-level method to disrupt memorized sequences at inference time. It operates entirely without access to the training corpus or model weights, addressing a key practical limitation of earlier techniques. By introducing this token-level intervention, the method effectively reduces both exact and near-verbatim generation while demonstrating easy integration with standard frameworks like HuggingFace, making it a highly scalable solution for production-level LLMs. This work highlights an emerging emphasis on practical, plug-and-play privacy interventions that can complement more computationally intensive training-time methods.
@@
Activation steering addresses memorization in LLMs by manipulating internal activations during inference [turner2023activation]. By injecting targeted perturbations into an LLM’s hidden states, activation steering can suppress or redirect the generation of memorized sequences while preserving overall model reasoning capabilities [suri2025mitigating]. Suri, et al. [suri2025mitigating] demonstrated that using a sparse autoencoder to identify activation patterns linked to memorized passages enables the construction of steering directions that reduced memorization by up to 60%, with minimal performance degradation. However, they note that steering effectiveness depends heavily on precise layer selection and steering strength.
@@
Understanding where memorization arises in LLMs is crucial to designing effective steering interventions. Stoehr, et al. [stoehr2024localizing] conducted a detailed mechanistic analysis of paragraph-length memorized sequences and found that memorization is often localized to specific LLM components. They showed that fine-tuning only a subset of the high-gradient weights was sufficient to erase memorized passages. Additionally, they identified a single attention head in an early layer that reliably activates in response to rare token combinations seen during training, triggering verbatim recall.
@@
Building on localization, Chang, et al. [chang2023localization] evaluated multiple methods for localizing memorized content by introducing two diagnostic benchmarks: injection, where known memorized sequences are inserted via fine-tuning a small set of weights, and deletion, which tests whether removing specific neurons erases a naturally memorized output. They compared several techniques, including activation-based heuristics [geva2022transformer], integrated gradients [dai2021knowledge], and pruning-based methods like Slimming and Hard Concrete [chang2023localization]. Their results showed that pruning-based approaches were most effective: for example, Hard Concrete was able to identify fewer than 0.5% of neurons whose removal led to a  60% drop in memorization accuracy. Both Stoehr, et al. [stoehr2024localizing] and Chang, et al. [chang2023localization] highlighted a key challenge: the neurons involved in one memory often contribute to others, complicating memory-specific interventions due to the risk of collateral forgetting.
@@
In Appendix’s Section 2.2, Table V provides a taxonomy of challenges in memorization mitigation by distinguishing between foundational scientific problems and the practical difficulties that arise at specific stages of the model lifecycle. The Foundational challenges, such as disentangling knowledge from regurgitation and verifying erasure, represent the core, cross-cutting scientific obstacles. The subsequent categories, Training-Time, Post-Training, and Inference-Time, detail how these fundamental issues manifest as concrete engineering trade-offs, such as the utility cost of differential privacy or the risk of catastrophic forgetting during unlearning. This structure offers a comprehensive overview, connecting the deep-rooted scientific difficulties to their practical consequences at each phase of model development and deployment.
@@ 6.4 Connections to Other Undesirable Behaviors
It remains open whether mitigating memorization also mitigates other undesirable behaviors of LLMs, such as hallucination, racial bias, and toxic content. For instance, does reducing memorization reduce hallucination on concepts related to memorized data? Is it possible that reducing the memorization of datapoints that include harmful racial stereotypes reduces the rates at which large models produce biased or toxic content? The researchers to date have not seen cross-cutting analysis examining whether reducing memorization can reduce other harmful behaviors in-tandem.
Mitigation efforts framed solely as the minimization of memorization overlook that an LLM’s utility is dependent on its capacity to store a vast repository of factual and linguistic patterns as parametric knowledge. The challenge is not suppression but disentanglement: separating the mechanisms that support beneficial, generalizable factual recall from those that cause the harmful regurgitation of instance-specific, non-generalizable data points. This perspective re-frames the objective from one of simple elimination to one of principled, selective control, motivating the critical open questions that follow on how to define, measure, and preserve this beneficial memorization while curbing the harmful.
@@
Open Questions 1. How can we scale up DP training, and at what cost and utility trade-off? 2. How can DP improve PEFT to reduce memorization? 3. How can we optimize activation steering for memorization removal? 4. How can post-training methods reduce memorization while preserving utility? 5. Can memorization mitigation methods be incorporated online during post-training, eg. by using memorization-detecting reward models during RLHF? 6. When does memorization contribute to versus harm utility or generalization? 7. Does mitigating memorization affect other undesirable behaviors, such as hallucination or generating toxic stereotypes?
OQ1. One direction for DP training is to develop content-sensitive DP, where the privacy budget ($\epsilon$) is allocated dynamically based on the semantic content. This reframes the challenge from applying uniform noise to intelligently learning to apply stricter protection for sensitive data while preserving utility on public knowledge, thus creating a more optimized and practical privacy-utility trade-off [li2021large, aghabagherloo2025impact, SHSP2024].
OQ2. While applying DP to PEFT methods is a practical first step, a formal theory for the composability of privacy guarantees when multiple private adapters are combined is critically lacking. Future work must therefore move beyond post-hoc DP application to designing novel PEFT architectures that are private-by-construction, ensuring mathematically provable guarantees for a secure, modular model ecosystem [chua2024mind, ma2024efficient, liu2024differentially].
OQ3. Current activation steering methods often focus on suppressing specific, known instances of memorization. A key open question is how to scale this from single-instance intervention to learning generalizable privacy policies at the activation level. Instead of merely blocking one secret, research should focus on identifying and modifying the underlying “reasoning paths” that lead to privacy violations, effectively teaching the model abstract rules like “do not reveal PII” that generalize [meng2022locating, suri2025mitigating, turner2023activation, stoehr2024localizing].
OQ4. As the machine unlearning paradigm shifts from one-shot erasures to a continuous governance framework, it creates critical challenges in developing scalable and formally verifiable deletion methods. Future work must therefore produce efficient and auditable techniques designed for the ongoing maintenance of LLMs as dynamic systems, capable of processing numerous requests without catastrophic forgetting [nguyen2025survey, cooper2024machine, chen2025parapoaligninglanguagemodels].
OQ5. The next evolution beyond a static memorization-detecting reward model is to create a dynamic adversarial alignment process. This would involve co-training a “red team” AI designed to continuously find and exploit memorization vulnerabilities, while the main LLM is rewarded for evading these adaptive attacks. This moves beyond fixed-preference optimization to a more robust, game-theoretic approach to achieving privacy during alignment [schulman2017proximal, rafailov2024directpreferenceoptimizationlanguage].
OQ6. The binary “good” versus “bad” memorization is insufficient; the future requires quantifying a spectrum of generalizability for all knowledge within an LLM. Research should develop information-theoretic or causal metrics to place any model output on a continuum from a non-generalizable, instance-specific regurgitation to a broadly applicable, foundational fact. This reframes the problem from a simple classification to a nuanced, risk-based assessment of knowledge provenance [feldman2020neural, hong2025reasoning, zhang2023counterfactual].
OQ7. Instead of tackling memorization, hallucination, and toxicity as separate problems, research should investigate a unified mechanistic hypothesis: that they are all symptoms of over-relying on spurious statistical correlations. Developing training methods rooted in causal representation learning could be a single, powerful intervention that simultaneously reduces harmful memorization, improves factuality, and curtails toxic stereotypes by forcing the model to learn the true data-generating process [bender2021dangers, siau2020artificial].
@@ 7 Privacy & legal risks of memorization
Memorization in LLMs poses serious security and privacy risks: memorized text can be directly leaked, undermining data confidentiality [smith2023identifying]. We discuss three major impact areas caused by memorization: personal data leakage, exposure of copyrighted or proprietary content, and broader legal and public consequences.
@@
Personal data leakage. As explored in [carlini2021extractingtrainingdatalarge, li2024llmpbeassessingdataprivacy, jagielski2023measuring, zhou2024quantifying], a significant risk for LLMs is leaking sensitive personal information. In [panda2024teach], researchers demonstrated a targeted neural phishing attack that achieved up to a 50% success rate in tricking a model into revealing PII during fine-tuning when injecting poisons during pretraining. Even when training data is intended to remain private, malicious attackers with the correct model access level can siphon memorized secrets [panda2024teach]. If an LLM reveals any PII, this could result in identity theft or other harms. Furthermore, for industries such as healthcare or customer service, where customers expect confidentiality, this would create significant liability for model operators under current privacy regulations.
@@
Copyright or proprietary content. Beyond PII, memorization in LLMs may also cause the reproduction of copyrighted or proprietary material such as books, code, etc. This turns the LLM into a potential conduit for the unauthorized distribution of protected content. As observed by [carlini2023quantifyingmemorizationneurallanguage], models may produce verbatim memorization from their training data. While we can only explicitly confirm what training data is used for open-source LLMs [gao2020pile, soldaini2024dolma], there do exist copyrighted books in EleutherAI’s PILE dataset, specifically, copyrighted books in its Books3 section. This is observed in [henderson2023foundation, zhang2022opt], where open models can regurgitate the first few pages of the Harry Potter books and “Oh, the Places You’ll Go!” by Dr. Seuss verbatim, raising concerns.
@@
Legal landscape. There are several ongoing cases that will shape the legal landscape of LLMs. The New York Times v. Microsoft Corp. lawsuit [nytimes_v_microsoft_2023] represents a pivotal legal challenge to generative AI and copyright, accompanied by similar cases, including Chabon v. OpenAI [chabon_v_openai_2023] where authors claiming copyright infringement in model training, and Doe v. GitHub [doe_v_github_2022] where plaintiffs allege verbatim code reproduction in violation of the Digital Millennium Copyright Act (DMCA). The Times claimed that OpenAI’s model trained on their copyrighted articles without proper license and that ChatGPT could regurgitate passages of those articles on request. In [nytimes_v_microsoft_2023], the publishers argue this behavior amounts to mass copyright infringement since the model’s outputs mimic and even compete with the Times’ content. High-profile copyright litigation, exemplified by [nytimes_v_microsoft_2023], has thrust LLM memorization from a technical concern into the public policy and business arenas, a shift underscored by empirical analyses of the lawsuit’s claims from privacy researchers [freeman2024exploringmemorizationcopyrightviolation].
In Appendix’s Section 2.3, Table VI provides a taxonomy that traces the path from core technical failures to their downstream legal and societal risks. It illustrates how specific Technical Challenges, such as the inability to guarantee verifiable erasure, directly create Legal & Regulatory Challenges, like the difficulty of complying with General Data Protection Regulation (GDPR)’s Right to Erasure, which in turn manifest as tangible Socio-Technical Challenges, such as the erosion of public trust. This structure is applied consistently across all impact areas, from personal data leakage to copyright infringement. The table’s central insight is that legal and societal risks are not independent but are fundamentally constrained by underlying technical limitations.
@@
OQ1. The next research frontier is to move beyond simple data filtering and train LLMs to be copyright-aware. This involves augmenting training corpora with copyright metadata and modifying the optimization objective to explicitly penalize the generation of text that is both verbatim and tagged as protected. This approach reframes copyright compliance from a static data-curation problem into a learnable, dynamic behavior, enabling the model to respect intellectual property as an intrinsic part of its generation process [vyas2023provable, karamolegkou2023copyright].
OQ2. Future memorization metrics must evolve from measuring superficial string similarity to quantifying the degree of transformative use, a cornerstone of fair use doctrine. The research challenge is to develop computational methods that assess not just what was reproduced, but how it was used—for example, distinguishing a derivative summary (potentially fair use) from a substitutive replication (potential infringement). This requires creating metrics that capture the functional and semantic relationship between an LLM’s output and a copyrighted source, providing a technical basis for legal arguments [henderson2023foundation, schwarzschild2024rethinking].
OQ3. Instead of searching for a single, universal threshold, the research community should develop a framework for context-aware memorization auditing. This framework would replace a simple percentage with a multi-faceted report assessing factors like the novelty of the memorized text content, its commercial value, and the “extraction effort” required by a user to elicit it. This re-orients the goal from finding a magic number to providing a rich, evidence-based dossier that legal experts can use to argue “substantial similarity” on a case-by-case basis [cooper2023report, nytimes_v_microsoft_2023].